16. At the time he purchased his variable annuity, Ahmed directed $5,000 of his premium into Subaccount A when the unit value was $10. A year later, the unit value had increased to $15. Assuming he made no additional premium deposits, what is the value of Ahmed's investment in Subaccount A now? a. $25,000 b. $16,666 c. $7,500 d. $8,333
Answer #1

Premium Invested = $5000

Investment Price = $10

Therefore, Units Purchased = Total Premium/Investment Price = 5000/10 = $500

Value after 1 year = $15

Therefore, Value of Investment afte 1 year = Number of units×Value after 1 year = 500×15 = $7500

Therefore, Correct Answer is (c)$7500.

(If this was helpful then please give postive rating to the answer. Thank You:)